# A Survey of AI Methods for Geometry Preparation and Mesh Generation in Engineering Simulation

### v1

@@ 4 Mesh Quality Prediction
Anticipating mesh quality before meshing is crucial for complex CAD models because early identification of regions likely to yield poor elements prevents costly iteration and focuses cleanup where it matters most. Instead of using meshing as a diagnostic that requires repeated cycles of mesh creation, quality inspection, control adjustments, and remeshing, an ML predictor can estimate local quality directly from B-rep features prior to element generation, surfacing hotspots in advance and guiding sizing and preparation choices. Prior work demonstrates this direction for tetrahedral meshing and positions the approach as algorithm-agnostic across meshers and parameter regimes [OwenIMR2019, OwenOSTI2022, OwenIMR2023]. Parry [Parry2025] presents an extensive B-rep-based study for mesh-quality prediction in defeaturing workflows, with emphasis on ensembles of decision trees and uncertainty estimation to support decision-making. In an earlier IMR research note, Parry compared a Random Forest against a CNN for local quality prediction, reporting better performance from the tree ensemble in this setting [ParryIMR2023].
@@
For each B-rep entity — vertex, curve (edge), and surface (face) — a fixed-length feature vector encodes local geometry and topology within a bounded neighborhood. Example features include curvatures, dihedral angles, proximity to neighboring features, valence, and length and area ratios. A training corpus of CAD parts is meshed under multiple target sizes and parameter settings to obtain localized quality labels near each entity, typically within one to two element edge lengths. Common labels include scaled Jacobian and inradius; definitions and implementations of these metrics are well established in the literature and libraries used by meshing tools [Knupp2000, Knupp2003, Verdict2007].
As with part classification, supervised learning is used: given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$, the goal is to learn a function $f$ such that $\hat{\mathbf{y}}=f(\mathbf{x})$. Ensembles of decision trees (Random Forests) [Breiman2001], implemented in scikit-learn [sklearn], deliver strong accuracy, fast inference, and interpretable feature importances on tabular B-rep features; Parry’s note finds RF outperforming a CNN baseline for local quality prediction [ParryIMR2023], consistent with results reported in applied workflows [OwenIMR2019, OwenOSTI2022, OwenIMR2023]. The mesh quality predictor is implemented in the Cubit® Geometry and Meshing Toolkit [CubitWebsite] as part of the defeaturing workflow, where it flags low-quality regions prior to meshing and informs subsequent corrective actions, reducing iteration and accelerating time to analysis [OwenIMR2019, OwenIMR2023, CubitWebsite]. Future work includes broader coverage of element types and quality metrics beyond tetrahedra, cross-mesher generalization, uncertainty-aware decision policies, and integration with graph- and B-rep-based neural models for joint geometry–topology reasoning.
@@ 5 Defeaturing
CAD models intended for manufacturing often contain small features such as holes, fillets, and engravings. These details are important for fabrication, but they rarely influence the physics of interest and can make meshing difficult by forcing refinement and producing poorly shaped elements. Defeaturing, the process of removing or simplifying such features, is therefore a critical step before simulation. The challenge is that analysts must decide which features can be removed safely, balancing geometry, physics, and meshing requirements. This makes defeaturing one of the most time-consuming and expertise-driven parts of analysis preparation.
Traditional approaches attempt to reduce this burden, but each has limitations. Tolerant meshing methods can bypass fine artifacts by relaxing geometric fidelity [ANSYS2025, MeshGemsCADSurf, Hu2018TetWild, Staten2024Morph], yet they risk discarding features important for fidelity. Manual tools such as CADfix [CADfix2025] and SpaceClaim [SpaceClaim2025] require significant analyst effort and deep experience. What is needed is a way to automate cleanup decisions that is both scalable and trustworthy.
@@
Two recent studies illustrate different directions for ML-driven defeaturing. Shinde et al. [Shinde2024] supervised directly on simulation outcomes for sheet metal parts, labeling operations by their effect on quantities of interest in finite element analysis. This produces labels aligned with engineering objectives, but it requires many expensive simulation runs and limits the approach to specific physics. In contrast, Owen et al. [OwenIMR2019] proposed a more general framework driven by mesh quality prediction. Their method first estimates baseline quality for each B-rep entity, as described in Sec. 4. Entities predicted to produce poor elements are flagged, and candidate defeaturing operations in Cubit® such as remove surface, composite surfaces, blunt tangency or collapse curve are then evaluated by separate ML models that predict the post-operation quality. The system ranks both entities and operations, presenting analysts with prioritized cleanup options.
@@
This capability has been released in Cubit® and is already in use, but it remains an active R&D effort. Current work is expanding the feature models to cover more operations, improving generalization, and incorporating reinforcement learning. Whereas the current supervised approach predicts static outcomes for single operations, reinforcement learning would enable multi-step procedures that carry out full defeaturing end-to-end. Another direction is incorporating analyst preferences by mining journal files that capture real sequences of cleanup commands. These data could be used to adapt operation rankings to physics context. For example, structural dynamics analysts may favor different cleanup operations than those working in structural mechanics. Together, these efforts point toward defeaturing that is not only automated but also adaptive and informed by both geometry and human expertise.
6 Unstructured Mesh Generation Methods
Defeaturing prepares CAD models for analysis by removing unnecessary complexity, but the next critical step is mesh generation itself. Even on simplified geometry, the robustness and quality of the mesh directly affect the accuracy, efficiency, and convergence of finite element analysis (FEA). Much of the prior discussion has assumed a CAD foundation, where geometry is defined by parametric solids and surfaces constructed in tools such as SolidWorks or Creo. However, many important applications do not begin with CAD at all. Medical imaging, reverse engineering, topology optimization, and scanned parts all provide geometry in implicit, voxel, or point-sampled form rather than as explicit CAD models. In such settings, the domain itself must first be constructed before meshing. This distinction motivates two categories of AI-enhanced approaches: those that construct the geometric domain from raw data, and those that generate unstructured meshes on an existing domain, whether CAD-based or reconstructed.
6.1 Geometry Construction
When CAD is absent, geometry must be built from discrete or implicit data before meshing is possible. Classical methods include Marching Cubes [lorensen1998marching], Dual Contouring, and level-set extraction, which generate watertight triangulated surfaces from voxelized data or distance fields. AI-driven variants improve robustness, feature preservation, and differentiability. DeepSDF and related implicit networks [park2019deepsdf, mescheder2019occupancy, peng2020convolutional, mildenhall2021nerf] represent shapes as continuous neural functions, from which smooth surfaces, normals, and curvature can be derived analytically [gropp2020implicit, larios2021deep, novello2022exploring]. Such neural implicits provide resolution independence and consistent geometry across scales, properties that can directly benefit downstream meshers by reducing boundary noise and preserving sharp features [dong2024neurcadrecon, dong2025neurcross].
Alternative formulations use neural parameterizations, predicting explicit spline or subdivision patches with direct CAD compatibility [sharma2020parsenet, williamson2025neural]. Compared to implicits, these models ease editing and boundary condition specification at the expense of handling complex topology. Hybrid approaches such as PoNQ [maruani2024ponq] and SparseFlex [he2025sparseflex] combine the flexibility of neural fields with the editability of meshes, enabling both high-resolution reconstruction and downstream editing. Differentiable variants of surface extraction, such as Deep Marching Cubes [liao2018deep] or learned Dual Contouring [chen2022neural, sundararaman2024self], relax discrete decisions into smooth surrogates, allowing gradients to flow through geometry reconstruction and making it possible to couple surface generation with optimization or learning objectives.
Although these methods originate largely in computer graphics, their relevance to simulation lies in providing well-conditioned geometric domains when CAD is not available. Applications include meshing of medical imaging data [zhang2016geometric, zhang2013challenges], topology-optimized structures, scanned parts, or any setting where geometry is defined implicitly. By producing smooth, consistent, and differentiable representations of geometry, they improve the reliability of subsequent unstructured mesh generation for FEA.
6.2 Meshing
Once a domain is available—either from CAD or reconstructed using the methods above—the challenge shifts to unstructured mesh generation itself. Classical algorithms such as Delaunay triangulation, advancing front, and paving remain central for tetrahedral, triangular, and quadrilateral meshing. These algorithms are mature and widely deployed, but still rely on heuristic rules for point insertion, front propagation, smoothing, and refinement. Such heuristics can be brittle across diverse geometries and often require manual intervention to achieve high-quality meshes.
AI offers the ability to augment these algorithms by learning adaptive policies that replace or refine heuristics. Reinforcement learning has been applied directly to advancing-front quad meshing: Pan et al. [pan2023reinforcement] formulate front updates as a sequential decision process and train a Soft Actor-Critic (SAC) policy that produces higher-quality meshes than rule-based strategies. Tong et al. [TONG2023102109] extend this with supervised pretraining and RL fine-tuning, improving singularity placement and cross-field alignment—two key challenges in quad meshing. More recently, Thacher, Persson, and Pan [Thacher2025] applied reinforcement learning with graph neural networks to generate then improve connectivity in triangular Delaunay mesh generation. Their method learns point placement and connectivity decisions that yield meshes of comparable or higher quality than established codes such as Triangle or DistMesh, while retaining the theoretical guarantees of Delaunay triangulation. This work demonstrates one of the clearest examples of AI directly enhancing a classical meshing algorithm used in FEA.
Complementary ideas from graphics illustrate how meshing kernels can be embedded into learning frameworks. Neural subdivision methods [liu2020neural, fruhauf2024quadify] retain classical combinatorics while learning residual vertex updates, improving smoothness and feature preservation. Mesh simplification guided by task objectives, such as MeshCNN [hanocka2019meshcnn], shows how decimation can be coupled to downstream goals, a principle analogous to adaptive refinement or coarsening in FEA. Hybrid pipelines, such as Quadify [fruhauf2024quadify], combine classical guarantees with neural refinement, providing a model for CAE workflows where correctness must be preserved but adaptivity is essential.
At present, AI-enhanced unstructured meshing remains at a low technology readiness level. Reinforcement learning for advancing front and Delaunay meshing are the most direct examples with FEA relevance, but they are still research prototypes. Neural implicit and differentiable geometry methods are more mature in graphics, yet their integration into CAE workflows is just beginning. Nonetheless, the potential impact is significant. By improving geometry conditioning in non-CAD domains and replacing brittle heuristics in meshing algorithms, AI can reduce manual intervention, increase automation, and enable tighter coupling between meshing and simulation objectives. Future work is likely to explore AI-guided point insertion in tetrahedral meshing, prediction of refinement and coarsening based on error estimators, and differentiable formulations that embed meshing within simulation-driven optimization loops. Together, these advances point toward a future where unstructured mesh generation evolves from a heuristic bottleneck into a data-driven, adaptive process integrated seamlessly with defeaturing, classification, segmentation, and downstream FEA.
7 Block Structured Meshing

### v2

@@ 4 Mesh Quality Prediction
Anticipating mesh quality before meshing can save substantial time for complex CAD models because early identification of regions likely to yield poor elements prevents costly trial and error and focuses cleanup where it matters most. Rather than using meshing as a diagnostic that requires repeated cycles of mesh creation, inspection, control adjustments, and remeshing, an ML predictor can estimate local quality directly from CAD boundary-representation (B-rep) features prior to element generation. In practice, the most useful output is a ranked list or heat map of likely problem regions, which allows an analyst to target sizing, defeaturing, or partitioning actions early [OwenIMR2019, OwenOSTI2022, OwenIMR2023].
Mesh quality metrics are widely used in this setting because they are inexpensive, broadly applicable, and correlate with common failure modes such as inverted or highly distorted elements [Knupp2000, Knupp2003, Verdict2007]. At the same time, mesh quality is only a proxy for what ultimately matters, namely the accuracy and reliability of simulation results. Learning simulation error or quantities of interest directly is often impractical because it requires solver outputs, boundary conditions, material models, and in many cases expensive reference solutions. A complementary line of work therefore predicts how model preparation choices affect downstream outcomes, such as estimating the impact of CAD defeaturing or simplification on heat transfer or CFD results [Danglade2014, Danglade2015]. More broadly, surrogate modeling efforts learn to predict solution fields or derived quantities directly on meshes [Pfaff2021MeshGraphNets], but these methods are typically problem-specific and place stronger demands on training data and generalization.
As with part classification, mesh quality prediction is naturally posed as supervised learning. Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$, the goal is to learn a function $f$ such that $\hat{\mathbf{y}}=f(\mathbf{x})$. The inputs $\mathbf{x}_{i}$ are constructed by assigning each B-rep entity (vertex, curve, and surface) a fixed-length feature vector that encodes local geometry and topology within a bounded neighborhood. Typical features include principal curvatures, dihedral angles, proximity to neighboring features, valence, and length and area ratios. This differs from part classification, where features are designed to characterize the part as a whole, often emphasizing global volume and overall shape. In mesh quality prediction, features are intentionally localized to the entity being evaluated and to the surrounding neighborhood that is most likely to influence the elements generated there.
@@
The labels $\mathbf{y}_{i}$ are obtained by meshing a corpus of CAD parts under multiple target sizes and meshing settings, then assigning localized quality values near each entity, typically within one to two element edge lengths. Common scalar targets include minimum scaled Jacobian and minimum scaled inner radius, with definitions and implementations available in the standard literature and in libraries used by meshing tools [Knupp2000, Knupp2003, Verdict2007].
With inputs and labels defined, ensembles of decision trees (Random Forests) [Breiman2001], implemented in scikit-learn [sklearn], provide an effective baseline for these tabular B-rep features because they deliver strong accuracy, fast inference, and interpretable feature importances. Parry compares a random forest against a CNN for local quality prediction and reports better performance from the tree ensemble in this setting [ParryIMR2023], consistent with results reported in applied workflows [OwenIMR2019, OwenOSTI2022, OwenIMR2023].
A secondary consideration, but an important one for tetrahedral meshing, is that quality outcomes can be inherently variable. Tetrahedral meshing is not space filling in the same way as structured approaches, and nondeterministic choices and small parameter changes can produce different local outcomes. This effect can be observed even on simple geometries: across repeated meshes of the same brick under random 3D rotations, the minimum scaled Jacobian for tets within two edge lengths of a surface varies by roughly 0.15. This variability motivates predicting a range, or the risk of failing a threshold, rather than a single deterministic score, and incorporating that uncertainty into how predictions are used [Parry2026-LearningTetQuality]. Practically, this is supported by repeated meshing across settings, followed by aggregation into distributional targets such as quantiles or threshold probabilities [Parry2026-LearningTetQuality].
Related efforts outside CAD-to-mesh pipelines also cast mesh quality assessment as a learning problem, including element-level predictors and graph-based evaluations learned from mesh connectivity [SpraveDrescher2021, Wang2022MeshQualityGNN]. These approaches are less central to CAD-driven pre-mesh decision making, but they reinforce the broader point that quality signals can be learned and can support automation.
The mesh quality predictor is implemented in the Cubit Geometry and Meshing Toolkit [CubitUserDoc1706_2025], where it flags low-quality regions prior to meshing and informs corrective actions that reduce iteration and accelerate time to analysis [OwenIMR2019, OwenIMR2023]. Future work includes broader coverage of element types and quality metrics beyond tetrahedra, improved cross-mesher generalization, uncertainty-aware decision policies, and tighter coupling between mesh-quality risk predictors and downstream estimators of simulation impact, including learned surrogates and goal-oriented error indicators.
Contributed by S. Owen
@@ 5 Defeaturing
CAD models intended for manufacturing often contain small features such as holes, fillets, embossed text, and fastener details. These features can be essential for fabrication, but they often have limited influence on the physics of interest and can make meshing expensive by forcing refinement and introducing poorly shaped elements. Defeaturing removes or simplifies such features so the mesh resolves what matters for analysis, without paying for unnecessary geometric detail.
Defeaturing in practice typically falls into three patterns. One option is to perform no defeaturing at all, allowing a conventional mesher to honor the full B-rep [MeshGemsCADSurf]. This can produce far too many elements and unnecessarily small elements, increasing compute cost with little gain in accuracy. A second option is to handle defeaturing implicitly during meshing by using a feature-size tolerance or geometric envelope so the mesh effectively ignores small details [Staten2024Morph, Hu2018TetWild]. This is usually fast and convenient, but it offers less control and can require iteration to confirm that important features were not suppressed. A third option is to explicitly edit the CAD model using geometry tools to remove or modify selected features. This offers the most control and often yields the best meshing outcome, but it is tedious and requires expertise. The latter two approaches are widely used in practice.
ML can reduce friction in tolerance-based meshing by predicting when it is likely to be safe, and it can accelerate explicit editing by identifying the entities most likely to create poor quality elements and recommending targeted CAD operations. In both cases, the goal is to reduce trial-and-error by estimating the expected meshing outcome, including the likely improvement, without requiring repeated meshing.
Both tolerant meshing and explicit CAD editing aim to reduce the defeaturing burden, but each comes with tradeoffs. Tolerant meshing can be an effective default in high-throughput workflows, and it is commonly available in commercial toolchains [ANSYS2025], but analysts still need safeguards to ensure that suppressed details do not change the physics of interest. Manual defeaturing tools such as CADfix [CADfix2025] and SpaceClaim [SpaceClaim2025] provide precise control, but they require significant analyst effort and deep experience with geometry operations and downstream meshing behavior. What is needed is a way to automate cleanup decisions that is both scalable and trustworthy, combining the speed of tolerant approaches with guidance and safeguards comparable to expert-driven editing.
@@
Two recent studies illustrate different directions for ML-driven defeaturing. Shinde et al. [Shinde2024] supervised directly on simulation outcomes for sheet metal parts, labeling operations by their effect on quantities of interest in finite element analysis. This produces labels aligned with engineering objectives, but it requires many expensive simulation runs and limits the approach to specific physics. In contrast, Owen et al. [OwenIMR2019] proposed a more general framework driven by mesh quality prediction. Their method first estimates baseline quality for each B-rep entity, as described in Sec. 4. Entities predicted to produce poor elements are flagged, and candidate defeaturing operations in Cubit [CubitUserDoc1706_2025] such as remove surface, composite surfaces, blunt tangency or collapse curve are then evaluated by separate ML models that predict the post-operation quality. The system ranks both entities and operations, presenting analysts with prioritized cleanup options.
@@
This capability has been released in Cubit [CubitUserDoc1706_2025] and is already in use, but it remains an active R&D effort. Current work is expanding the feature models to cover more operations, improving generalization, and incorporating reinforcement learning. Whereas the current supervised approach predicts static outcomes for single operations, reinforcement learning would enable multi-step procedures that carry out full defeaturing end-to-end. Another direction is incorporating analyst preferences by mining journal files that capture real sequences of cleanup commands. These data could be used to adapt operation rankings to physics context. For example, structural dynamics analysts may favor different cleanup operations than those working in structural mechanics. Together, these efforts point toward defeaturing that is not only automated but also adaptive and informed by both geometry and human expertise.
Contributed by S. Owen
6 Geometry Construction from Implicit Data
Defeaturing prepares CAD models for analysis by removing unnecessary complexity, but many simulation workflows do not begin with CAD. Medical imaging, reverse engineering, topology optimization, and scanned parts often provide geometry as voxels, implicit fields, or point samples rather than as parametric solids and surfaces authored in CAD systems. In these settings, the first meshing problem is not discretization, but domain construction. The objective is to recover a well-conditioned geometric description that is watertight, faithful to important features, and stable under refinement so that downstream unstructured meshing and finite element analysis (FEA) do not inherit artifacts from the input representation. These non-CAD inputs are especially common in medical image-based modeling, where robust domain construction is a prerequisite for reliable analysis meshes [zhang2016geometric, zhang2013challenges].
Classical pipelines convert implicit or discretized data into a boundary representation using methods such as Marching Cubes [lorensen1998marching], Dual Contouring [Zhang3DMeshing2005, ZhangHexMeshing2006], and level-set extraction. These approaches are reliable workhorses, but their outputs can be sensitive to resolution, noise, and thresholding choices. For simulation, that sensitivity often appears as boundary roughness, inconsistent normals, and fragile sharp features, all of which complicate sizing, boundary-layer construction, and mesh quality control.
The methods reviewed here fall along three complementary axes: (i) continuous representations that decouple geometric fidelity from any single tessellation, (ii) differentiable reformulations of discrete extraction so reconstruction can be guided by downstream losses, and (iii) learned module replacement and end-to-end designs that retain trusted structure while improving robustness and adaptivity.
Continuous representations. Neural implicit models represent a shape as the zero level set of a learned field $f_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}$, with signed distance functions (SDFs), occupancy fields, and radiance-field variants being common choices [park2019deepsdf, mescheder2019occupancy, peng2020convolutional, mildenhall2021nerf]. Practical training strategies such as Eikonal regularization encourage well-behaved distance fields and improve conditioning [gropp2020implicit]. A key benefit for simulation pipelines is that continuous fields support differentiable geometric queries. Using automatic differentiation, normals and curvature-related quantities can be derived from $\nabla f_{\theta}$ and $\nabla^{2}f_{\theta}$ [goldman2005curvature, larios2021deep, novello2022exploring]. This supports feature-aware reconstruction that is more stable across discretizations, and it has been used in downstream workflows such as CAD reconstruction with improved sharp-feature retention and more robust cross-field generation for quad meshing [dong2024neurcadrecon, dong2025neurcross]. Continuous formulations can also encode distance structure on surfaces, for example neural geodesic fields that learn distances and paths, with decoupled variants improving scalability while maintaining accuracy in their reported settings [zhang2023neurogf, chencoupled].
Neural parameterizations provide a complementary continuous description by mapping low-dimensional coordinates to explicit surfaces such as NURBS, subdivision surfaces, or spline patches, giving stronger topological control and closer alignment with CAD workflows. ParSeNet predicts parametric control structures from raw 3D data to fit surfaces robustly [sharma2020parsenet]. Spherical Neural Surfaces provide a continuous formulation that stabilizes differential operators across irregular meshes and varying discretizations, improving the stability of spectral computations in that framework [williamson2025neural]. In addition, hybrid representations seek a practical compromise between flexibility and editability. PoNQ uses a learned quadric-error style representation to compress and reconstruct editable meshes efficiently [maruani2024ponq], while SparseFlex leverages sparse neural fields to capture fine detail and thin structures under rendering or multi-view photometric supervision [he2025sparseflex].
Differentiable reformulations of extraction. Even when geometry is represented continuously, most CAE pipelines ultimately require an explicit surface for boundary-conforming meshing. This motivates differentiable reformulations of classical extraction steps so that losses tied to geometry quality can influence reconstruction. Marching Cubes is non-differentiable because of discrete case tables and topology updates [lorensen1998marching]. Deep Marching Cubes relaxes thresholding and case selection into learnable surrogates, enabling supervision with geometric losses such as normal consistency while still producing a mesh output [liao2018deep].
Building on this idea, related work explores alternative differentiable isosurfacing formulations that better align extraction with optimization and downstream objectives. Deep Marching Tetrahedra (DMTet) uses a tetrahedral discretization and differentiable marching tetrahedra to extract high-resolution triangle meshes from learned fields [shen2021dmtet]. FlexiCubes targets stability and mesh quality during gradient-based optimization by introducing a flexible isosurface extractor that improves robustness when extraction is embedded in learning or inverse-design loops [shen2023flexicubes]. Dual Contouring can likewise benefit from learned vertex placement and self-supervised geometric consistency, reducing dependence on ground-truth meshes while improving sharp-feature preservation [chen2022neural, sundararaman2024self]. These relaxations provide usable gradients, but practical systems must still control instability in local topology decisions to avoid frequent flips and inconsistent surfaces [liao2018deep, chen2022neural, sundararaman2024self].
Learned modules and end-to-end designs. A third axis replaces brittle or hand-tuned components with learned modules, while still producing explicit geometric outputs and preserving the overall pipeline structure. Learned vertex placement in Dual Contouring is a representative example, where the neural component replaces heuristic optimization and improves sharp-feature preservation under realistic input variability [chen2022neural, sundararaman2024self]. Hybrid designs such as SparseFlex similarly couple a neural representation to multi-view photometric constraints, allowing reconstruction quality to be driven by dense supervision signals without requiring ground-truth meshes [he2025sparseflex]. For CAE, the appeal of these designs is pragmatic: they retain familiar stages (reconstruction, extraction, meshing) while reducing manual cleanup and improving robustness across acquisition modalities.
In a complementary direction, recent generative models produce explicit surface meshes directly. Meshtron [hao2024meshtron] demonstrates high-fidelity triangle mesh synthesis at scale, which is best viewed in this context as a geometry representation step. For simulation workflows, such outputs may still require conditioning, validation, or remeshing to meet analysis constraints.
From a technology readiness perspective, geometry construction from implicit data is comparatively mature in computer graphics, but its translation into routine CAE workflows remains at an early stage. The opportunity is clear: better-conditioned domains reduce manual cleanup, improve meshing robustness, and make it easier to connect reconstruction to simulation objectives. Near-term progress is likely to focus on topology control and feature-aware validation tailored to analysis needs, together with tighter integration into downstream sizing, defeaturing, and mesh-quality prediction.
Contributed by N. Lei, X. Gu
7 Unstructured Mesh Generation
Once a domain is available, either from CAD or constructed from implicit data as described above, the workflow bottleneck shifts to unstructured mesh generation. Mesh robustness and element quality directly influence accuracy, efficiency, and convergence in FEA. Classical unstructured meshing algorithms are mature and widely deployed, but they rely on heuristic decisions that can be brittle across geometry classes and often require manual intervention to achieve reliable quality.
Delaunay-based methods, advancing-front techniques, and paving remain central for tetrahedral, triangular, and quadrilateral meshing. These methods combine proven geometric structure with a sequence of design choices, including point insertion, front updates, smoothing, and local refinement. In practice, it is these choices, rather than the high-level algorithm, that limit robustness. Small changes in geometry, sizing fields, or boundary complexity can degrade element quality or trigger topological failures, and the resulting fixes are often manual and time consuming.
The main opportunities for learning arise in the parts of unstructured meshing where heuristic choices control success or failure. Heuristic rules are typically designed for common cases, whereas production geometries are diverse and increasingly originate from non-CAD sources. Learned policies can adapt to local geometric context while still operating within classical algorithmic frameworks, preserving structural guarantees while replacing brittle decision logic with data-driven components.
One representative line of work formalizes meshing as a sequential decision process. Advancing-front all-quad meshing can be posed as reinforcement learning, where a Soft Actor-Critic policy learns front updates that generalize across boundary conditions and sizing fields [pan2023reinforcement]. Building on this idea, SRL-Assisted AFM combines supervised pretraining with reinforcement-learning fine-tuning to control singularities, improving cross-field alignment and local element quality on challenging planar domains [tong2023srl]. In practice, these methods target robust quad layouts in multiply connected regions by making singularity placement an explicit part of the learned control policy, which helps maintain consistency of the front progression while preserving element quality near holes and other constraints.
Related work also applies learning to Delaunay-style triangular meshing, using graph-based policies to guide point placement and connectivity while preserving the underlying Delaunay structure [Thacher2025]. In parallel, several efforts embed learning into established pipelines rather than replacing them end-to-end. Neural subdivision variants retain Loop or Catmull–Clark combinatorics and learn residual vertex updates conditioned on local geometry [liu2020neural], which can improve fairness and feature preservation during refinement. Hybrid pipelines also demonstrate how classical kernels can be paired with learned refinement while keeping trusted structure intact. Quadify is a representative example that integrates differentiable Catmull–Clark refinement into a pipeline that builds on field-aligned meshing foundations such as Instant Field-Aligned Mesh [fruhauf2024quadify, jakob2015instant]. Task-driven simplification provides another instructive pattern: MeshCNN learns edge importance so that decimation preserves structures relevant to downstream objectives [hanocka2019meshcnn]. In simulation settings, the analogous opportunity is learned refinement and coarsening guided by solution features and error indicators, with the learned component acting as an adaptive policy inside a verified meshing framework.
AI-enhanced unstructured meshing remains at a low technology readiness level. Most approaches are research prototypes, and practical adoption must address verification, robustness across geometry distributions, and compatibility with CAE requirements such as boundary layers, sizing constraints, and quality guarantees. The potential impact is significant because reducing manual intervention in meshing directly lowers analysis turnaround time and makes larger design spaces tractable. Promising directions include AI-guided point insertion and refinement in tetrahedral meshing, learned refinement and coarsening informed by error estimation, and tighter coupling between meshing decisions and simulation feedback. Over time, these advances can help move unstructured meshing from a heuristic bottleneck toward a more adaptive, objective-aware component that integrates smoothly with defeaturing, segmentation, and downstream FEA.
Contributed by N. Lei, X. Gu
8 Structured Meshing